I would like to know how to solve this problem:
Find the value of x using the formula x=50-6y when y=4.
Please show me how to this and explain it.
Thank you for your assistance.

Answer by shawtybabi09(1)

Well first locate the equation. The equation is x= 50 -6*y. *=times or multiplication. Well you already know what y is y = 4. So you take y which is 4 and plug into the equation or sub-situate for y. making the equation x = 50 - 6*4
First you would do the multiplication because the order of operations are:
P- ( )parenthesis
E- exponents
M- multiplication
A- addition
S- subtraction
6*4= 24
so now the equation is x= 50 - 24
do the subtraction, 50 - 24 = 26 and there is your answer
x=26